OpenStudy (anonymous):

Number of words which can be formed using all the letter of the word HIP HIP HURRAY in which all H's lie somewhere between R's is

OpenStudy (anonymous):

@rational

OpenStudy (rational):

There are `12` letters in total ``` H : 3 I : 2 P : 2 U : 1 R : 2 A : 1 Y : 1 ```

OpenStudy (rational):

Put this string of 5 letters in a bag : `RHHHR`

OpenStudy (rational):

then we have \(12-5+1 = 9\) objects to arrange, out of which I and P repeat 2 times : \[\frac{9!}{2!2!}\]

OpenStudy (anonymous):

read the question again

OpenStudy (rational):

Oops! the H's need not be adjacent okay

OpenStudy (anonymous):

yep..i thought of making cases but that would be silly... and I did a question in the past and i do have a bare hint of that arranging some letters and rest the letters will be arranged by themselves or something like that

OpenStudy (anonymous):

leave that question... got a new one If a number is selected at random from the set of all five digit numbers in which sum of the digits is equal to 43, then find the probability that this number is divisible by 11 is

OpenStudy (anonymous):

@perl

OpenStudy (thomas5267):

\[ a=\sum_{k=0}^4a_k10^k\equiv0\pmod{11}\iff\sum_{k=0}^4(-1)^ka_k\equiv0\pmod{11} \] {9,9,9,9,7}, {9,9,9,8,8} are the only possible combinations of digits such that they sum to 43. \[9-7+9-9+9=11\equiv0\pmod{11}\] And my iPad is running out of battery so I will leave the rest for you

OpenStudy (anonymous):

um quite confused!

OpenStudy (anonymous):

a number is divisible by 11 when alternate digits sum is equal and idk how to arrange 99988

OpenStudy (thomas5267):

For {9,9,9,8,8}, there are only three situations. Either two 8's are negative, one 8 is positive and one 8 is negative or two 8's are positve. So sum of digits are either 3*9-2*8=11, 9-9+9-8+8=9 or 8-9+8-9+9=7. For {9,9,9,9,7}, 9-7+9-9+9=11 is only solution.

OpenStudy (thomas5267):

So 98989 is the only number that is divisible by 11 formed by the set of digit {9,9,9,8,8} as 9-8+9-8+9=11. The only valid exchange of digits such that the number will still be divisible by 11 is switching the two 8's, which will result in the same number. For {9,9,9,9,7}, 97999 and 99979 are both valid solution as 9-7+9-9+9=9-9+9-7+9=11. So there are three solutions. You should be able to figure out the rest.

OpenStudy (anonymous):

one more doubt... :D how to examine if a number is divisible by 11? what i have in mind is alternate number sum should be equal like 121; 1+1 =2 so I dont get what do you mean by two 8's are negative

OpenStudy (thomas5267):

By two 8's are negative, I mean that in 9-8+9-8+9 the two 8's are subtracted from other numbers.

OpenStudy (anonymous):

@thomas5267 got any idea for the original permutation question?

OpenStudy (thomas5267):

Do the words formed have to be valid in English or is this just a combinatorial question?

OpenStudy (thomas5267):

H : 3 R : 2 I : 2 P : 2 U : 1 A : 1 Y : 1 Consider the case where R are at both ends of the word. Then there are \(\dfrac{10!}{3!2!2!1!1!1!}=151200\) possibilities. I postulate that the combinations are equal to\[ \frac{7!}{2!\,2!}\sum_{k=0}^7\binom{n+3}{3}(8-n)=997920 \]

OpenStudy (thomas5267):

Let me explain the divisible by 11 question. Part a, the set of digits such that the sum is 43. The maximum sum of digit of a 5 digit number is 45. \[ 9+9+9+9+9=45\\ 9+9+9+9+9-2=43\\ 9+9+9+9+7=43\\ 9+9+9+(9-1)+(9-1)=43\\ 9+9+9+8+8=43 \] So the two sets are {9,9,9,8,8} and {9,9,9,9,7}. Part b, numbers that are divisible for 11 constructed using the two sets of digits. \[ a=\sum_{k=0}^4a_k10^k\\ a=\sum_{k=0}^4a_k10^k\equiv0\pmod{11}\iff\sum_{k=0}^4a_k(-1)^k\equiv0\pmod{11}\\ a_0-a_1+a_2-a_3+a_4\equiv0\pmod{11} \] Consider the set {9,9,9,8,8}. We have to add three elements of the set and subtract two elements of the set such that the result is 11n (as \(a_0-a_1+a_2-a_3+a_4\equiv0\pmod{11}\)). Either we add two 8's, we add one 8 and subtract one 8, or we subtract two 8's. 8+8-9+9-9=7 (add two 8's) 9-9+9-8+8=9 (add one 8 and subtract one 8) 9+9+9-8-8=11 (subtract two 8's) So 98989 is the only number constructed using {9,9,9,8,8} as digits that is divisible by 11. (We have subtracted two 8's, so it must be in \(a_1\) and \(a_3\). Swapping the two 8's still yields 98989.) Consider the set {9,9,9,9,7}. Again, add three elements and subtract two such that the results equal 11n. This time, we either add 7 or subtract 7. 9+9+7-9-9=7 (add 7) 9+9+9-9-7=11 (subtract 7) So 7 is either \(a_1\) or \(a_3\) as we have subtracted 7. 99979 (\(7=a_1\)) and 97999 (\(7=a_3\)) are the two numbers. There are 3 numbers that are divisible by 11 in the set of five digits numbers such that the sum of digits = 43. The probability of finding such number in the set is left as an exercise for the reader XD.